OpenStudy (anonymous):

let n be positive integer that \[\log_{2} \log_{2} \log_{2} \log_{2} \log_{2} n<0<\log_{2} \log_{2} \log_{2} \log_{2}n\] Let l be the number of digits in the binary expansion of n. Then the minimum and the maximum possible values of l are

OpenStudy (anonymous):

@aroub ? @bibby ? @campbell_st ?

OpenStudy (anonymous):

@Hero ?

OpenStudy (anonymous):

@nincompoop ?

OpenStudy (anonymous):

@sammixboo ?

OpenStudy (aum):

\[\large \log_{2} \log_{2} \log_{2} \log_{2} \log_{2} n<0<\log_{2} \log_{2} \log_{2} \log_{2}n \\ \large \text{If } a \lt b \lt c, \text{ then }2^a < 2^b < 2^c. \normalsize ~~\text{Therefore, } \\ \large 2^{\log_{2} \log_{2} \log_{2} \log_{2} \log_{2} n} < 2^0 < 2^{\log_{2} \log_{2} \log_{2} \log_{2} n} \normalsize\text{ which simplifies to one less log:}\\ \large \log_{2} \log_{2} \log_{2} \log_{2} n<1<\log_{2} \log_{2} \log_{2}n ~~~~ \normalsize \text{ since }~2^{\log_2(x)} = x\\ \text{Raise them again to the power of 2 and simplify. There will be one less log:} \\ \large \log_{2} \log_{2} \log_{2} n<2^1=2< \log_{2} \log_{2}n \\ \text{Raise them again to the power of 2 and simplify. } \\ \large \log_{2} \log_{2} n<2^2=4< \log_{2}n \\ \text{Raise them again to the power of 2 and simplify. } \\ \large \log_{2} n<2^4=16< n. ~~~ \\ \text{n has to be > 16. This is the lower limit of 'n'} \\ \text{Raise them again to the power of 2 and simplify. } \\ \large n < 2^{16} = 65536 < 2^n \\ \text{n has to be < 65536. This is the upper limit of 'n'} \\ \]Since n is an integer, the lowest value is 17 and the highest value is 65535.

OpenStudy (aum):

17 in binary is: 1 0001 (5 binary digits) 65535 in binary is: 1111 1111 1111 1111 (16 binary digits)